### v2

# AdaAugment: A Tuning-Free and Adaptive Approach to Enhance Data Augmentation

Abstract
Data augmentation (DA) is widely employed to improve the generalization performance of deep models. However, most existing DA methods use augmentation operations with random magnitudes throughout training. While this fosters diversity, it can also inevitably introduce uncontrolled variability in augmented data, which may cause misalignment with the evolving training status of the target models. Both theoretical and empirical findings suggest that this misalignment increases the risks of underfitting and overfitting. To address these limitations, we propose AdaAugment, an innovative and tuning-free Adaptive Augmentation method that utilizes reinforcement learning to dynamically adjust augmentation magnitudes for individual training samples based on real-time feedback from the target network. Specifically, AdaAugment features a dual-model architecture consisting of a policy network and a target network, which are jointly optimized to effectively adapt augmentation magnitudes. The policy network optimizes the variability within the augmented data, while the target network utilizes the adaptively augmented samples for training. Extensive experiments across benchmark datasets and deep architectures demonstrate that AdaAugment consistently outperforms other state-of-the-art DA methods in effectiveness while maintaining remarkable efficiency.
Data Augmentation
1 Introduction
(a) Traditional DA
(b) Adaptive DA
Figure 1: Comparison between traditional DA methods and adaptive DA methods.
Data augmentation (DA) is a crucial and highly effective technique to alleviate overfitting in the training of deep neural networks (Xu et al., 2023; Yang et al., 2022), resulting in more generalized and reliable models. However, it is important to note that existing DA methods primarily rely on the utilization of totally random or manually devised augmentation magnitudes during training, which can lead to suboptimal training scenarios and require more significant engineering efforts of human experts to tune the hyper-parameters given a new dataset. For instance, information deletion-based DA methods (DeVries & Taylor, 2017; Chen et al., 2020; Yang et al., 2023) randomly erase some sub-regions in images to create augmented samples, leading to random augmentation strengths. Additionally, automatic DA approaches (Cubuk et al., 2020; Müller & Hutter, 2021; Lin et al., 2023a) often entail significant computational overhead to optimize augmentation strategies for each specific dataset before the target model training. The magnitudes of these augmentations are predetermined without any adaptation throughout the training process. Consequently, these methods may fail to capture the unique characteristics of each dataset and the evolving state of network training.
In fact, using augmentation magnitudes introduces inherent randomness, which can increase data diversity but also introduces uncontrolled variability into the augmented data. Such randomness in data variability may not optimally align with the evolving training status of deep models, thereby introducing severe side effects to the training process. For instance, during the initial stage of training, when models generally exhibit weaker generalization capabilities, substantial data variability can induce noise and distributional shifts, potentially leading to underfitting phenomena (Soviany et al., 2022). In contrast, in the later stages of training, limited data variability may elevate the risk of overfitting (Yang et al., 2023). Consequently, this misalignment in the current DA methodologies increases the risks of both underfitting and overfitting, which can ultimately have adverse effects on the generalization performance of models (Gong et al., 2021; Yang et al., 2024). This is a challenge often overlooked in the field.
To tackle these challenges, it is crucial to adopt an adaptive modulation of augmented data variability based on real-time feedback from the target models. As illustrated in Fig. 1, we present a comparison between traditional DA and adaptive DA mechanisms. In contrast to conventional approaches that rely on random or predefined augmentation magnitudes, adaptive DA methods dynamically adjust the magnitudes of DA operations based on real-time feedback from the target network. Notably, adaptive DA focuses on optimizing augmentation strengths rather than the specifics of the operations themselves. This adaptive strategy offers an effective countermeasure to mitigate the aforementioned challenges, thereby unleashing the full potential of DA and enhancing the generalization capabilities of deep models. One straightforward solution to achieve adaptive DA is to create a formal measure that indicates the real-time learning status of each training sample. This measure can then be used to determine the appropriate level of augmentation strengths. Nevertheless, theoretical analyses have identified that determining the overfitting or underfitting risks of a learning algorithm is undecidable (Bashir et al., 2020; Sehra et al., 2021; Li et al., 2024). Consequently, formulating a definitive measure of learning status for training samples remains a formidable challenge.
To address the limitations and challenges above, in this work, we propose AdaAugment, an innovative and tuning-free Adaptive data Augmentation method. AdaAugment leverages a reinforcement learning algorithm to adaptively determine specific augmentation magnitudes for each training sample without relying on any hand-crafted measures. Central to AdaAugment is a dual-model architecture: the policy network and the target network. The policy network learns the policy that determines the magnitudes of augmentation operations for each training sample based on the real-time feedback from the target network during training. The target network simultaneously utilizes these adaptively augmented samples for training. Both networks are jointly optimized, eliminating the need for separate re-training of the target network and enhancing the practical feasibility of our approach. The learned policy adjusts the variability within the augmented data to align with the learning status of the target model, thereby optimizing the introduced variability into the augmented data. Specifically, in the training of the policy network, we estimate the risks of underfitting and overfitting by deriving losses from fully augmented and non-augmented data, respectively. These two losses are then compared with those derived from data adaptively augmented by AdaAugment, serving as reward signals. Extensive experimental results across various benchmark datasets demonstrate the superior performance of AdaAugment compared to state-of-the-art (SOTA) DA methods. Furthermore, complexity analyses of AdaAugment verify that it merely imposes minimal parameters and computational overhead, highlighting its high efficiency. Therefore, AdaAugment achieves a commendable balance between effectiveness and efficiency, delivering high performance without introducing excessive computational complexity. For example, AdaAugment can achieve more than 1% improvements over other SOTA DA methods when training on Tiny-ImageNet without any architectural modifications or additional regularization. In summary, we highlight our contributions as follows:
  1. 1.
    We propose AdaAugment, an innovative and tuning-free adaptive DA approach that leverages reinforcement learning techniques to dynamically adapt augmentation magnitudes for individual training samples based on real-time feedback from the target network.
  2. 2.
    AdaAugment features a dual model architecture that jointly optimizes a policy network and a target network. This facilitates concurrent policy formulation and enhances task performance effectively.
  3. 3.
    To the best of our knowledge, this is the first work that identifies that the adaptive DA mechanism can dynamically improve models’ generalization.
  4. 4.
    Extensive experiments conducted on multiple benchmark datasets and deep networks show that AdaAugment outperforms existing SOTA DA methods. Importantly, AdaAugment achieves this superior performance while incurring minimal additional training overhead, thus ensuring competitive efficiency.
2 Related Work
2.1 Data Augmentation
Data augmentation has played a crucial role in enhancing the generalization capabilities of deep neural networks. Typically, these methods have primarily relied on augmentations characterized by random or predefined magnitudes to introduce diversity into the training data.
Among these methods, Cutout (DeVries & Taylor, 2017) is one of the most widely used techniques, which randomly masks out one or more square regions within training images. Random Erasing (Zhong et al., 2020) randomly selects a rectangular region within an image and erases its pixels with random values. Similarly, Hide-and-Seek (HaS) randomly hides patches in a training image, improving object localization accuracy and the generalization ability of deep models. GridMask (Chen et al., 2020) employs structured dropping regions within the input images. Since these methods may easily introduce noise and ambiguity into the augmented data, AdvMask (Yang et al., 2023) is proposed to identify classification-discriminative information in images and structurally drop some sub-regions that contain critical points for augmentation. Meanwhile, Mixup (Zhang et al., 2018) and CutMix (Yun et al., 2019) mix random information from two or more images to synthesize the augmented data. However, these DA approaches primarily focus on data transformations and often overlook the training status of models. This oversight makes it challenging to alleviate actual overfitting and underfitting risks accordingly through online adjustments to augmentation strengths.
Figure 2: The general framework of the proposed AdaAugment.
Automatic DA methods, such as AutoAugment (Cubuk et al., 2019), Fast-AutoAugment (Lim et al., 2019), and RandAugment (Cubuk et al., 2020), leverage reinforcement learning or grid search to search existing policies for the optimal combination of DA operations on different image datasets in an offline manner. Similarly, Adversarial AutoAugment (Zhang et al., 2019) utilizes fixed augmentation space and rewards the policy yielding the lowest accuracy, causing the policy distribution to shift towards progressively stronger augmentations throughout training. Meanwhile, TrivialAugment (Müller & Hutter, 2021) employs the same augmentation space obtained by these automatic DA methods and applies only a single augmentation operation to each image during training. MetaAugment (Rajendran et al., 2020) utilizes a static augmentation space, and its augmentation policy network outputs weights for augmented data loss. However, the magnitudes of augmentation policies employed in these methods remain fixed or are randomly sampled during online training, leading to the uncontrollable extent of data transformations. SelectAugment (Lin et al., 2023a) employs hierarchical reinforcement learning to acquire online policies that determine the portion of the augmented data and whether each individual sample should be augmented. However, since the specific augmentation transformation applied to each sample is AutoAugment, Mixup, or CutMix, the DA strengths employed are still uncontrollable. KeepAugment (Gong et al., 2021) is proposed to detect and preserve the salient regions of the images during augmentation.
The work (Fawzi et al., 2016) proposes an adaptive DA method that focuses on identifying small transformations that yield maximal classification loss on the augmented data. The work (Lee et al., 2020) leverages impact modeling via the influence function to learn differentiable DA transformation. The work (Xu & Zhao, 2022) randomly determines the types and magnitudes of DA operations for batch-level data and updates DA’s parameters along the gradient direction of the target network loss. The work (Zhang et al., 2023) proposes AdDA for contrastive learning, which allows the network to adjust the augmentation compositions and achieve more generalizable representations. Adaaug (Cheung & Yeung, 2021) learns class-dependent and potentially instance-dependent augmentations via a differentiable workflow. TeachAugment (Suzuki, 2022) transforms data so that they are adversarial to the target model. However, previous adaptive DA methods typically sample augmentation policies with fixed magnitudes and mainly aim to alleviate overfitting risks. More recently, the work (Yang et al., 2024) evaluates DA’s efficacy using similarity and diversity metrics, revealing their varying importance across datasets, suggesting the merit of adaptively adjusting augmentation strengths for optimal balance. Therefore, in contrast to existing DA methods, AdaAugment estimates both underfitting and overfitting risks during training and adaptively adjusts augmentation magnitudes to reduce both risks during online training.
2.2 Reinforcement Learning
Reinforcement learning (RL) is to learn a sequence of actions within an interactive environment by trial and error that maximizes the expected reward (Le et al., 2022; Li, 2023), which obtains a wide range of applications, such as autonomous driving (Kiran et al., 2022) and recommender systems (Lin et al., 2023b), etc. In the realm of RL algorithms, there are two fundamental categories: value optimization and policy optimization methods. Value optimization methods primarily revolve around the estimation of an optimal value function, which subsequently serves as a basis for deriving the optimal policy (Byeon, 2023). Conversely, policy optimization methods estimate the optimal behavior policy without estimating the value functions (Sutton et al., 1999). Moreover, the actor-critic framework, widely adopted in RL, combines the strengths of both value-based and policy-based RL methods (Shakya et al., 2023). This framework encompasses two components: the actor, which is responsible for learning the policy function, and the critic, which is tasked with evaluating the actor’s chosen actions by estimating the value function. This dual mechanism ensures more stable and efficient learning (Grondman et al., 2012), such as Advantage Actor-Critic (A2C) (Mnih et al., 2016).
3 AdaAugment
Overview
The primary objective of AdaAugment is to alleviate underfitting and overfitting risks through an adaptive adjustment of DA strengths during training. This adaptive adjustment can be formalized as a decision-making problem on a per-sample basis, thereby eliminating the need for manually crafted metrics. Fig. 2 illustrates the dual-model framework of AdaAugment: While training the target network, a policy network is simultaneously introduced to dynamically optimize the magnitudes of DA operations in the augmentation process. This dual-model framework jointly optimizes two networks, obviating the necessity for separate re-training of the target network and enabling real-time adjustments of augmentation strengths based on the training progress of the target network. Specifically, the policy network optimizes the magnitudes at the sample level in the current training epoch. Then, in the next epoch, the data augmentation process applies augmentation operations to the training samples with these corresponding magnitudes. The adaptively augmented samples are then utilized for optimizing target network training.
Preliminary
An RL task can be formalized as a Markov Decision Process (MDP), which comprises the following components: a state space $\mathcal{S}$, an action space $\mathcal{A}$, a transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ that represents the probability of transitioning from one state to another by taking actions, a reward function $\mathcal{R}$, a discount factor $\gamma\in[0,1]$, and a time step $T$. Given a state $\boldsymbol{s}\in\mathcal{S}$, an RL agent determines an action $a\in\mathcal{A}$ with the policy $\pi(a\mid\boldsymbol{s})$. With these elements, the objective of an RL task is to find an optimal policy, denoted as $\pi^{*}$, that maximizes the expected cumulative reward within the given MDP framework. Meanwhile, suppose that the training dataset $\mathcal{D}$ consists of $N$ training samples, each of the form $(\boldsymbol{x},\boldsymbol{y})\in\mathcal{D}$. Here, $\boldsymbol{x}$ denotes the original data, and $\boldsymbol{y}$ is a $k$-dimensional vector of 0s and 1s indicating the true label of $\boldsymbol{x}$, where $k$ is the total number of classes. We define the augmentation operation as $e(m,\boldsymbol{x})$, where $e$ is drawn from an augmentation space $\mathcal{E}$, and $m$ corresponds to the magnitude of $e$. Compared to previous works (Cubuk et al., 2020; Müller & Hutter, 2021), our approach adaptively determines the magnitude in the augmentation space $\mathcal{E}$ rather than assigning a predefined value before training.
State Design
Since the objective of RL is to dynamically determine the appropriate magnitudes for each sample, the state $\boldsymbol{s}\in\mathcal{S}$ should consider three factors: the inherent difficulty associated with each sample, the current training status of the models (e.g., feature extraction capabilities), and the intensity of augmentation operations w.r.t. the former two factors. This multifaceted consideration of state variables is essential for effective RL. Notably, the feature map plays a vital role in this process by providing feedback from the model (Huang et al., 2023; Zhang et al., 2021), encapsulating both the inherent difficulty of the sample and the model’s real-time feature extraction capability. To illustrate this, as shown in Fig. 2, the state vector $\boldsymbol{s}$ in AdaAugment encodes the feature maps of both the non-augmented sample $\boldsymbol{x}$ and adaptively augmented sample $\tilde{\boldsymbol{x}}$, denoted as $\boldsymbol{s}_{none}$ and $\boldsymbol{s}_{ada}$, respectively. Here, $\boldsymbol{s}_{none}$ captures the inherent difficulty of sample $\boldsymbol{x}$ w.r.t. the current target network training status. Meanwhile, $\boldsymbol{s}_{ada}$ encodes the augmentation operation strength applied to $\tilde{\boldsymbol{x}}$ w.r.t. both the real-time target network status and the inherent sample difficulty. By utilizing this information, AdaAugment effectively aligns data variability with the evolving training status of the target networks.
Action Design
The policy determines the augmentation magnitudes $\boldsymbol{m}$ for the augmented data. Despite the randomness in the composition of each mini-batch during training, the magnitudes $\boldsymbol{m}$ operate on a per-sample basis, corresponding to each training sample. For simplicity, we denote the magnitudes of the current mini-batch data as $\boldsymbol{m}$, where the dimension of $\boldsymbol{m}$ equals the batch size, and each $m\in\boldsymbol{m}$ is strictly bounded within the interval $[0,1]$. When $m=0$, no augmentation is applied, while $m=1$ indicates the maximum magnitudes for the corresponding augmentation operations. In this way, magnitudes closer to 0 yield samples more similar to the original, whereas magnitudes closer to 1 produce more diverse data. The adaptively augmented sample $\tilde{\boldsymbol{x}}$ can be obtained by:
$\tilde{\boldsymbol{x}}=e(m,\boldsymbol{x}),$ (1)
where $e$ is a random augmentation operation and $m$ is determined based on the action policy. Therefore, Eq. (1) enables us to optimize the similarity-diversity preference of the augmented data. More importantly, this adjustment on the similarity-diversity preference reflects AdaAugment’s effectiveness in alleviating overfitting and underfitting risks.
Reward Function
Consider a target classification model $f_{\theta}$, parameterized by $\theta$, an input sample $\boldsymbol{x}\in\mathbb{R}^{n}$, and $f_{\theta}(\boldsymbol{x})$ denotes the network output. Let $\mathcal{L}$ denote the cross-entropy loss, and $\mathcal{L}(f_{\theta}(\boldsymbol{x}),\boldsymbol{y})$ denote the loss item for the original sample $(\boldsymbol{x},\boldsymbol{y})$. Our proposed method aims to mitigate overfitting and underfitting risks by controlling the magnitudes of data augmentation operations based on feedback from the target model. To this end, we define three loss items based on our augmentation strategy. Firstly, $\mathcal{L}_{full}(f_{\theta}(\boldsymbol{x}^{+}),\boldsymbol{y})$ represents the loss of sample $\boldsymbol{x}$ with the maximal magnitude, i.e., $\boldsymbol{x}^{+}=e(m=1,\boldsymbol{x})$. Secondly, $\mathcal{L}_{none}(f_{\theta}(\boldsymbol{x}^{-}),\boldsymbol{y})$ denotes the loss for the non-augmented sample, i.e., $\boldsymbol{x}^{-}=e(m=0,\boldsymbol{x})=\boldsymbol{x}$. Lastly, based on Eq. (1), the loss for the adaptively augmented data is denoted as $\mathcal{L}_{ada}(f_{\theta}(\tilde{\boldsymbol{x}},\boldsymbol{y}))$, with the augmentation magnitude determined by the actor network. Inspired by the curriculum learning (Soviany et al., 2022), we formulate the reward function as follows:
$r=\lambda(\mathcal{L}_{full}-\mathcal{L}_{ada})+(1-\lambda)(\mathcal{L}_{ada}-\mathcal{L}_{none}),$ (2)
where $\lambda\in[0,1]$ is an adjusting factor, which is initialized at 1 and gradually decreases to 0 during the training process.
Policy Learning
The policy aims to determine the instance-level magnitudes for augmentation operations. For policy learning, we utilize the widely-used A2C algorithm (Mnih et al., 2016), which consists of an actor network $\theta_{a}$ and a critic network $\theta_{c}$, as depicted in Fig. 2. The actor network learns the policy, i.e., the probability distribution over actions given a particular state, $\pi(a\mid\boldsymbol{s})$. Meanwhile, the critic network aims to estimate the value associated with a particular state, denoted as $V^{\pi}(\boldsymbol{s})$.
To update the actor and critic networks, we reformulate the loss functions tailored to our specific problem scenario. Specifically, the loss function for updating $\theta_{a}$ is defined as:
$\mathcal{L}_{\text{actor}}=-\log\pi_{\theta_{a}}(a\mid s_{none})(r+\gamma V_{\theta_{c}}(s_{ada})-V_{\theta_{c}}(s_{none})).$ (3)
Meanwhile, the loss function for updating $\theta_{c}$ is defined as:
$\mathcal{L}_{\text{critic}}=\mathbb{E}[(r+\gamma V_{\theta_{c}}(s_{ada})-V_{\theta_{c}}(s_{none}))^{2}].$ (4)
To provide a comprehensive insight into AdaAugment, Algorithm 1 offers a detailed algorithmic procedure.
Algorithm 1 The general workflow of AdaAugment.
0:  dataset $\mathcal{D}$, batch size $B$, an augmentation space $\mathcal{E}$, cross-entropy loss $\mathcal{L}$, total number of epochs $T$
1:  Randomly initialize $f_{\theta}$, $\theta_{a}$, and $\theta_{c}$
2:  $\boldsymbol{m}\leftarrow\boldsymbol{0}$
3:  for $t=0$: $T-1$ do
4:     Sample a mini-batch $\{\boldsymbol{x}_{i},y_{i}\}_{i=1}^{B}$ from $\mathcal{D}$
5:     Sample a random augmentation operation $e$ from $\mathcal{E}$
6:     Generate the adaptively augmented sample $\tilde{\boldsymbol{x}}_{i}$ with corresponding magnitude $m\in\boldsymbol{m}$ and fully augmented sample $\boldsymbol{x}_{i}^{+}$ according to Eq. (1)
7:     Compute the loss items $\mathcal{L}_{none}(f_{\theta}(\boldsymbol{x}_{i}),\boldsymbol{y}_{i})$, $\mathcal{L}_{full}(f_{\theta}(\boldsymbol{x}^{+}_{i}),\boldsymbol{y}_{i})$, and $\mathcal{L}_{ada}(f_{\theta}(\tilde{\boldsymbol{x}}_{i},\boldsymbol{y}_{i}))$
8:     Determine the corresponding adaptive magnitudes $\boldsymbol{m}$ using $\theta_{a}$
9:     Calculate reward value according to Eq. (2)
10:     Update $\theta_{a}$ and $\theta_{c}$ according to Eq. (3) and Eq. (4)
11:     Update $f_{\theta}$ based on $\mathcal{L}_{ada}(f_{\theta}(\tilde{\boldsymbol{x}}_{i},\boldsymbol{y}_{i}))$
12:  end for
12:  $f_{\theta}$
Theoretical Analysis
In the realm of data augmentation, the augmentation magnitude is approximately proportional to the loss values of the augmented samples, i.e., $m\propto\mathcal{L}$. Therefore, the following inequality approximately holds: $\mathcal{L}_{none}\leq\mathcal{L}_{ada}\leq\mathcal{L}_{full}$, which suggests that $\mathcal{L}_{none}$ and $\mathcal{L}_{full}$ can serve as indicators of potential overfitting and underfitting risks associated with augmented samples, respectively. Meanwhile, these loss values continuously evolve as the training progresses.
Regarding the reward function in Eq. (2), the term $\mathcal{L}_{full}-\mathcal{L}_{ada}$ holds significant importance during the initial training phases. Consequently, the policy network tends to employ relatively modest augmentation magnitudes, aiming to maximize the disparity between $\mathcal{L}_{full}$ and $\mathcal{L}_{ada}$. Therefore, more similar augmented samples can encourage models to capture broader features or patterns, accelerating rapid convergence during early training stages (Liu & Mirzasoleiman, 2022; Hou et al., 2023). As training advances, the focus shifts towards $\mathcal{L}_{ada}-\mathcal{L}_{none}$, prompting the policy network to apply greater augmentation magnitudes, trying to increase the difference between $\mathcal{L}_{ada}$ and $\mathcal{L}_{none}$. This yields more diverse augmented samples, mitigating overfitting risks. Thus, the methodology aligns with the curriculum learning principle (Soviany et al., 2022; Azad et al., 2023), wherein models initially learn patterns from similar samples and benefit from diverse samples in later training stages.
Additionally, to investigate the similarity-diversity trade-off of AdaAugment throughout the training process, as depicted in Fig. 3, we evaluate the similarity and diversity measures (Yang et al., 2024) at various training stages. Specifically, we calculate the average magnitude values for all samples at the intervals of 50 epochs. Notably, it can be observed that the diversity measure increases and the similarity measure decreases gradually as the training progresses, which aligns with our theoretical analysis. This highlights AdaAugment’s inherent adaptability in adjusting DA strengths to harmonize with the evolving training dynamics of the target models.
Figure 3: Evolving magnitudes throughout the training process.
Table 1: Test accuracy (%) on CIFAR-10/100 (average $\pm$ std). * means results reported in previous papers.
Dataset Method ResNet-18 ResNet-50 WRN-28-10 ShakeShake
CIFAR-10 baseline 95.28$\pm$0.14* 95.66$\pm$0.08* 95.52$\pm$0.11* 94.90$\pm$0.07*
CutMix 96.64$\pm$0.62* 96.81$\pm$0.10* 96.93$\pm$0.10* 96.47$\pm$0.07
Cutout 96.01$\pm$0.18* 95.81$\pm$0.17 96.92$\pm$0.09 96.96$\pm$0.09*
HaS 96.10$\pm$0.14* 95.60$\pm$0.15 96.94$\pm$0.08 96.89$\pm$0.10*
GridMask 96.38$\pm$0.17 96.15$\pm$0.19 97.23$\pm$0.09 96.91$\pm$0.12
RandomErasing 95.69$\pm$0.10 95.82$\pm$0.17 96.92$\pm$0.09 96.46$\pm$0.13*
AdvMask 96.44$\pm$0.15* 96.69$\pm$0.10* 97.02$\pm$0.05* 97.03$\pm$0.12*
AutoAugment 96.51$\pm$0.10* 96.59$\pm$0.04* 96.99$\pm$0.06 97.30$\pm$0.11
FAA 95.99$\pm$0.13 96.69$\pm$0.16 97.30$\pm$0.24 96.42$\pm$0.12
RandAugment 96.47$\pm$0.32 96.25$\pm$0.06 96.94$\pm$0.13* 97.05$\pm$0.15
DADA 95.58$\pm$0.06 95.61$\pm$0.14 97.30$\pm$0.13* 97.30$\pm$0.14*
TrivialAugment 96.28$\pm$0.10 97.07$\pm$0.08 97.18$\pm$0.11 97.30$\pm$0.10
Ours 96.75$\pm$0.06 97.34$\pm$0.13 97.66$\pm$0.07 97.41$\pm$0.06
CIFAR-100 baseline 77.54$\pm$0.19* 77.41$\pm$0.27* 78.96$\pm$0.25* 76.65$\pm$0.14*
CutMix 79.45$\pm$0.17 81.24$\pm$0.14 82.67$\pm$0.22 79.57$\pm$0.10
Cutout 78.04$\pm$0.10* 78.62$\pm$0.25 79.84$\pm$0.14 77.37$\pm$0.28
HaS 78.19$\pm$0.23 78.76$\pm$0.24 80.22$\pm$0.16 76.89$\pm$0.33
GridMask 75.23$\pm$0.21 78.38$\pm$0.22 80.40$\pm$0.20 77.28$\pm$0.38
AdvMask 78.43$\pm$0.18* 78.99$\pm$0.31* 80.70$\pm$0.25* 79.96$\pm$0.27*
RandomErasing 75.97$\pm$0.11* 77.79$\pm$0.32 80.57$\pm$0.15 77.30$\pm$0.18
AutoAugment 79.38$\pm$0.20 81.34$\pm$0.29 82.21$\pm$0.17 82.19$\pm$0.19
FAA 79.11$\pm$0.09 79.08$\pm$0.12 79.95$\pm$0.12 81.39$\pm$0.16
RandAugment 78.30$\pm$0.15 80.95$\pm$0.22 82.90$\pm$0.29* 80.00$\pm$0.29
DADA 78.28$\pm$0.22 80.25$\pm$0.28 82.50$\pm$0.26* 80.98$\pm$0.15
TrivialAugment 78.67$\pm$0.19 81.34$\pm$0.18 82.75$\pm$0.26 82.14$\pm$0.16
Ours 79.84$\pm$0.27 81.46$\pm$0.12 83.23$\pm$0.23 82.82$\pm$0.25
4 Experiments
Extensive experiments are conducted to evaluate the performance of our proposed method.
Experimental Settings
To the best of our knowledge, image classification is the most widely used task for evaluating the effectiveness of DA methods (Lin et al., 2023a). We evaluate the effectiveness of our method across several benchmark datasets, including CIFAR10/100 (Krizhevsky & Hinton, 2010) and Tiny-ImageNet (Chrabaszcz et al., 2017). Additionally, we employ transfer learning to evaluate the efficacy of our method in enhancing model generalization performance. Moreover, we conduct effect visualization on MNIST (LeCun et al., 1998) and perform a convergence analysis of AdaAugment throughout the training process. A complexity analysis is incorporated to verify the computational efficiency and resource utilization of AdaAugment. Meanwhile, the effectiveness of AdaAugment on long-tail biased ImageNet-LT and Places-LT (Liu et al., 2019) is provided in Appendix LABEL:appendix:oltr.
Comparison with State-of-the-arts
We compare our method with 10 most representative and commonly-used data augmentation methods, including CutMix (Yun et al., 2019), Cutout (DeVries & Taylor, 2017), HaS (Singh & Lee, 2017), GridMask (Chen et al., 2020), AdvMask (Yang et al., 2023), RandomErasing (Zhong et al., 2020), AutoAugment (Cubuk et al., 2019), Fast-AutoAugment (Lim et al., 2019), RandAugment (Cubuk et al., 2020), and TrivialAugment (Müller & Hutter, 2021).
Implementation Details
The discount factor $\gamma$ in Eq. (3) and Eq. (4) is set as 0.99 in our method. We closely follow previous works (Yang et al., 2023; Müller & Hutter, 2021) in our setup. Specifically, all images are preprocessed by dividing each pixel value by 255 and normalized by the dataset statistics. For CIFAR-10 and CIFAR-100, we use ResNet architectures, including ResNet-18/50 (He et al., 2016), Wide-ResNet-28-10 (WRN-28-10) (Zagoruyko & Komodakis, 2016), and Shake-Shake-26x32 (Gastaldi, 2017). We train 1800 epochs with cosine learning rate decay for Shake-Shake using SGD with Nesterov Momentum and a learning rate of 0.01, a batch size of 256, $1e^{-3}$ weight decay. We train 300 epochs for all other networks using SGD with Nesterov Momentum and a learning rate of 0.1, a batch size of 128, a $5e^{-4}$ weight decay, and cosine learning rate decay. For Tiny-ImageNet, we resize the images into 64$\times$64, initialize models with ImageNet pre-trained weights, and then fine-tune models using various augmentations. The baseline approach involves solely padding and horizontal flipping. For a fair comparison, all methods are implemented with the same training configurations, conducting experiments across three independent random trials.
4.1 Results on CIFAR-10 and CIFAR-100
In Table 1, we evaluate the efficacy of AdaAugment on CIFAR-10 and CIFAR-100, employing various widely-used deep networks, including ResNet-18/50, Wide-ResNet-28-10, and ShakeShake-26-32. It can be observed that AdaAugment consistently demonstrates superior performance in improving the accuracy of these networks when compared to existing SOTA DA methods. Notably, our method demonstrates significant improvements for both datasets. For example, AdaAugment achieves improvements on the baseline model’s performance of ResNet-18, ResNet-50, Wide-ResNet-28-10, and ShakeShake-26-32 by 1.47%, 1.66%, 2.14%, and 2.51%, respectively. This superior performance of AdaAugment can be attributed to its adaptive adjustment of augmentation magnitudes based on the real-time feedback received from the target network during training. This adjustment effectively mitigates the risks of underfitting and overfitting, resulting in improved generalization capabilities.
4.2 Results on Tiny-ImageNet
In this section, we evaluate the efficacy of AdaAugment on the Tiny-ImageNet dataset using ResNet-18/50, Wide-ResNet-50-2 (WRN-50-2), and ResNext50 (Xie et al., 2017), as summarized in Table 2. It can be observed that AdaAugment demonstrates a significant enhancement in classification accuracy across various architectures, outperforming other widely employed DA methods by a large margin. Particularly, AdaAugment elevates the accuracy by 9.87%, 5.5%, 1.52%, and 1.16% on ResNet-18, ResNet-50, WRN-50-2, and ResNext-50, respectively. This performance enhancement underscores the efficacy of our novel adaptive augmentation mechanism. Thus, we demonstrate the superior efficacy and potential of AdaAugment for enhancing model performance.
Table 2: Image classification accuracy (%) on Tiny-ImageNet dataset is summarized in this table (average $\pm$ std).
Method ResNet-18 ResNet-50 WRN-50-2 ResNext-50
baseline 61.38$\pm$0.99 73.61$\pm$0.43 81.55$\pm$1.24 79.76$\pm$1.89
CutMix 64.09$\pm$0.30 76.41$\pm$0.27 82.32$\pm$0.46 81.31$\pm$1.00
Cutout 68.67$\pm$1.06 77.45$\pm$0.42 82.27$\pm$1.55 81.16$\pm$0.78
HaS 63.51$\pm$0.58 75.32$\pm$0.59 81.77$\pm$1.16 80.52$\pm$1.88
GridMask 62.72$\pm$0.91 77.88$\pm$2.50 82.25$\pm$1.47 81.05$\pm$1.33
AdvMask 65.29$\pm$0.20 78.84$\pm$0.28 82.87$\pm$0.55 81.38$\pm$1.54
RandomErasing 64.00$\pm$0.37 75.33$\pm$1.58 81.89$\pm$1.40 81.52$\pm$1.68
AutoAugment 67.28$\pm$1.40 75.29$\pm$2.40 79.99$\pm$2.20 81.28$\pm$0.33
FAA 68.15$\pm$0.70 75.11$\pm$2.70 82.90$\pm$0.92 81.04$\pm$1.92
RandAugment 65.67$\pm$1.10 75.87$\pm$1.76 82.25$\pm$1.02 80.36$\pm$0.62
DADA 70.03$\pm$0.10 78.61$\pm$0.34 83.03$\pm$0.18 81.15$\pm$0.34
TrivialAugment 69.97$\pm$0.96 78.41$\pm$0.39 82.16$\pm$0.32 80.91$\pm$2.26
Ours 71.25$\pm$0.64 79.11$\pm$1.51 83.07$\pm$0.78 81.92$\pm$0.29
4.3 Results on ImageNet
Table 3: Top-1 accuracy (%) on ImageNet dataset.
baseline HaS Cutout CutMix GM Mixup AA FAA RA DADA TA AdaAugment
77.1 77.2 77.1 77.2 77.9 77.0 77.6 77.6 77.6 77.5 77.9 78.2
We evaluate AdaAugment on large-scale IamgeNet dataset (Krizhevsky et al., 2017).
4.4 Transfer Learning
Transfer learning plays a pivotal role in assessing the models’ transferability (Zhuang et al., 2021), a common practice in data augmentation (Gong et al., 2021; Yang et al., 2023). To quantitatively assess the efficacy of various DA methods, we assess the transferred accuracy of models using different DA methods. Specifically, we pre-train ResNet-50 models on the CIFAR-100 dataset utilizing different augmentations, followed by a subsequent fine-tuning of these models on the CIFAR-10 dataset. Thus, superior DA techniques lead to higher transferred test accuracy.
In Fig. 4, we show the transferred accuracy achieved by various DA methods. While the discrepancies in transferred accuracy may appear modest, it is noteworthy that AdaAugment consistently outperforms others. This consistent and effective performance underscores our method’s capability to enhance model transferability performance.
Figure 4: Transferred test accuracy of various DA methods.
4.5 Analytical Results
4.5.1 Visualization of Effectiveness
Since the objective of DA is to enhance the generalization performance of models, in this section, we conduct a comparative analysis of effectiveness visualization between models utilizing AdaAugment and those without. Specifically, we leverage the t-SNE algorithm to visualize the feature embedding of the MNIST test set, utilizing feature maps derived from different trained models. Theoretically, models that showcase more robust generalization capabilities tend to excel in feature extraction ability. Fig. 5 presents the t-SNE visualization results for the baseline model and the model trained with AdaAugment. Compared to Fig. 5(a), it can be seen that Fig. 5(b) exhibits an optimized geometric structure characterized by enhanced inter-cluster separation and intra-cluster compactness.
Furthermore, to furnish a quantitative evaluation of these embedding results, we employ the Dunn Index (DI) (Ncir et al., 2021) as an evaluative metric. The DI is mathematically defined as follows:
$DI=\frac{\min_{1\leq i\neq j\leq m}\delta\left(C_{i},C_{j}\right)}{\max_{1\leq j\leq m}\Delta_{j}}.$ (5)
Here, separation $\delta\left(C_{i},C_{j}\right)$ denotes the inter-cluster distance metric between clusters $C_{i}$ and $C_{j}$, and compactness $\Delta_{j}$ calculates the mean distance between all pairs within each cluster. A higher DI value signifies superior clustering. In Fig. 5, the DI values for the baseline and AdaAugment are $5.03\times 10^{-5}$ and $1.05\times 10^{-4}$, respectively. Consequently, AdaAugment achieves a DI value that is 108.7% higher than the baseline, demonstrating its efficacy in enhancing model performance. Based on both quantitative and qualitative analysis, we validate the substantial effectiveness of AdaAugment in improving model performance.
(a) baseline
DI=5.03×$\times 10^{-5}$
(b) AdaAugment
DI=$1.1\times 10^{-4}$
Figure 5: Visualization of effectiveness using t-SNE algorithm. DI: Dunn index.
4.5.2 Complexity Analysis
Table 4: The additional architecture parameters and training time of the auxiliary policy network. We report average $\pm$ std.
Models FLOPs Parameters GPU hours Improved Acc.
ResNet-18 1.82G +0.15M +0.43h±$\pm$0.03 +1.47%$\pm$0.06
WRN-28-10 5.25G +0.19M +0.41h±$\pm$0.03 +2.04%$\pm$0.07
Most existing DA methods typically rely on predefined random magnitude values during training. Although this mechanism incurs minimal additional computational overhead in online training, it may not be optimally aligned with the dynamic evolution of deep models, potentially introducing undesirable side effects. To address this, AdaAugment incorporates an auxiliary policy network into online training to determine the magnitudes of DA operations dynamically. Although this adjustment inevitably introduces slightly additional parameters and computational complexity, it is a deliberate trade-off to achieve more effective DA.
In this section, we thoroughly analyze the parameter and time complexity of AdaAugment to assess its efficiency. Since the policy network’s parameter complexity relates to the target network’s feature space, we report the parameter complexity on two different deep networks, i.e., ResNet-18 and WRN-28-10. In Table 4, we present the complexity analysis for AdaAugment using the CIFAR-10 dataset, following the methodology employed in prior studies (Cubuk et al., 2019; Lim et al., 2019; Lin et al., 2023a). It can be observed that the policy network’s parameter complexity results in a mere 1.3% increase for ResNet-18 (with 11.7M parameters) and a 0.52% increase for WRN-28-10 (with 36.5M parameters). Furthermore, the overall training costs experience an increase of less than 0.5 GPU hours for these networks. These findings highlight the negligible training overhead of AdaAugment.
Notably, despite the marginal increases in training costs, AdaAugment delivers substantial accuracy improvements, achieving a notable enhancement of 1.47% and 2.04% for ResNet-18 and WRN-28-10, respectively. This improvement underscores AdaAugment’s remarkable effectiveness-efficiency trade-off, demonstrating its capacity to enhance model performance while incurring minimal training costs.
4.5.3 Convergence Analysis
In this section, we conduct a convergence analysis of AdaAugment throughout the entire training process and compare it against two of the most representative DA methods, AutoAugment and TrivialAugment. Specifically, to present the dynamic evolution of test errors during training, we train ResNet-18 models on CIFAR-10 using a multi-step learning rate decay schedule. The learning rate is initialized as 0.1 and multiplied by 0.2 at epochs 60, 120, 160, 220, and 280. From Fig. 6, it can be observed that AdaAugment significantly enhances model performance, particularly after the second learning rate drop. Moreover, even after the initial learning rate drop, AdaAugment consistently maintains lower error rates than other methods. These empirical findings not only underscore the superior efficacy of AdaAugment in terms of model performance enhancement but also emphasize its capacity to expedite the convergence of the model during training.
Figure 6: Convergence analysis of training process.
5 Conclusion
In this work, we propose AdaAugment, an innovative and tuning-free method for adaptive data augmentation. By adaptively adjusting the magnitudes of DA operations based on the real-time feedback obtained during the training of the target model, AdaAugment optimizes the data variability. This dynamic adjustment ultimately mitigates the risks associated with both underfitting and overfitting, thereby optimizing model training and enhancing model generalization capabilities. Through extensive empirical evaluations, we demonstrate that our adaptive augmentation mechanism helps to significantly improve the performance across various benchmark datasets with competitive efficiency. In the future, we hope to generalize AdaAugment to more tasks.
6 Impact Statements
The method proposed in this paper, grounded in both theory and experimentation, aims to advance research in data augmentation within the field of deep learning. The experimental data is sourced entirely from standard datasets, eliminating any ethical concerns. Additionally, there are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.